A solution freezes at -2.30°C. At what temperature will it boil?
1
Expert's answer
2019-09-16T04:06:02-0400

The сryoscopic constant for water is Kf =1.853 K·kg/mol.

The ebullioscopic constant for water is Kb = 0.512 K·kg/mol.

The molality of the solution b = ∆Tf / Kf = 2.30K / 1.853 K·kg/mol = 1.2412 mol/kg.

∆Tb = b × Kb = 1.2412mol/kg × 0.512 K·kg/mol = 0.6355 K ≈ 0.64K.

Therefore, it will boil at 100.64⁰C.